# Forge: Foundational Optimization Representations from Graph Embeddings

### v1

@@ 5.2.1 Numerical Results: Integrality Gap Prediction
Test Instances & Setup: We start with the pre-trained Forge model, fine-tuned with a prediction head using the labeled data, and then predict the integrality gap of 50 ‘very-hard’ instances each of Combinatorial Auction (CA), Set Cover (SC), Generalized Independent Set (GIS), and Minimum Vertex Cover (MVC). Given the prediction, a pseudo-cut is generated by adjusting the initial LP relaxation objective, specifically, by increasing it for minimization problems (or decreasing it for maximization problems). This pseudo-cut is incorporated into the original formulation as an additional constraint, enforcing that the integral objective must exceed (or fall below) the computed pseudo-cut value.
Prediction Accuracy: We start with the accuracy of our predictions on the test instances. We measure the deviation in terms of mean absolute error between the true integrality gap and the gap predicted by Forge. On the ‘very-hard’ instances, Forge achieves a deviation of $15.42\%$, $13.55\%$, $12.03\%$ and $19.077\%$ for CA, SC, GIS and MVC instances respectively. Note that MVC instances were not part of the train set.
@@
Comparison with the Commercial MIP Solver: We compare the default Gurobi solver on these ‘very-hard‘ instances with and without our predicted pseudo-cut. Figure 5 shows the primal gap averaged over 50 instances of ‘very-hard’ instances for these test problems with a time limit of 3600 seconds. Without an exception, across all problem types, the use of pseudo-cuts generated by Forge embeddings consistently results in significantly lower primal gaps over time. The gains range from 30% to 85%. The default solver attains immediate improvements early on in the search and our pseudo-cuts make these gains more pronounced immediately. Recall that this only needed 50 instances from each problem type.
@@
Comparison with SotA ML Methods: To further evaluate generalization across problem types and sizes, we compare against the setup described by  Li et al. (2025), where a GNN is trained on 38,256 instances from 643 generated problem types and tested on 11,584 instances spanning 157 problem types.
We use Forge as is, not using any additional training instances apart from 450 D-MIPLIB instances, and test on 17,500 previously unseen instances spanning 400 generated problem types. Forge achieves a mean deviation of 18.63% in integrality gap prediction, outperforming the reported 20.14% deviation by almost 10%.
@@ 5.3 Warm-Start Prediction & Guiding The Search
To address this issue, we produce 5 feasible solutions for a given instance. Variables from the instance are grouped based on how many of the 5 solutions each variable was repeated in. A variable is considered a “negative” variable if and only if it appears in none of the 5 solutions. Next, we use triplet loss (Schroff et al., 2015) to fine tune the embeddings of the variables, where the triplet loss is given by
@@
Here, $a$ is the ‘anchor’ node, $p$ is the ‘positive’ node, $n$ is the ‘negative’ node and $d$ is a distance function (euclidean distance in our case). The goal of triplet loss is to minimize the distance between the ‘anchor’ and ‘positive’ nodes while ensuring that the ‘negative’ node is at least ‘margin’ distance away from the ‘anchor’ node (margin is set to 2 in our case). This is shown in Figure 4(C).
In our setting, all variables appearing in the same number of solutions are treated as ‘positive’ and ‘anchor’ pairs. A key challenge in using triplet loss is to find good negative nodes, as picking nodes that are trivially negative does not aid learning. To pick negative nodes, we pick variables that have not appeared in any solution but are closest to the positive node in the unsupervised embedding space. Finally, the Forge model is fine-tuned to predict warm start variables using a combination of triplet loss and BCE loss.
Training Instances: We collect 100 instances each from Combinatorial Auction (easy, medium), Set Cover (easy, medium, hard) and Generalized Independent Set (easy, medium) for a total of 700 training instances.
Label Collection: To generate training labels, we solve a given instance using Gurobi and a solution pool of 5. Here, we set a larger time limit of 300 seconds since we would like to get closer to the optimal solution.
@@ 5.3.1 Numerical Results: Warm-Start Prediction
Figure 7: Performance comparison of three variants of PS-Gurobi: the original model, a version augmented with 64-dimensional random vectors, and a version augmented with Forge embeddings (reduced to 64 dimensions via PCA). Evaluation is conducted on 50 instances each of CA and GIS problems. The Forge-augmented PS-Gurobi consistently outperforms the baseline, achieving average reductions in primal gap of 41.07% for CA and 50.51% for GISP.
@@
Since Forge embeddings for each node are 1024-dimensional, we apply Principal Component Analysis (PCA) to reduce the dimensionality to 64 dimensions, thereby minimizing architectural modifications to the original model. To further ensure that any observed performance gains are attributed to the semantic content of the Forge embeddings rather than increased model capacity, we also conduct a control experiment by augmenting PS-Gurobi with 64-dimensional random vectors. This comparison allows us to isolate the contribution of the learned embeddings to overall performance. We evaluate this setup on the same 50 instances each of the CA and GIS problems - two problem types that are common to both Forge and PS-Gurobi. Results are shown in Figure 7 where PS-Gurobi + Forge consistently outperforms the standalone PS-Gurobi with the gain in primal gaps being 12.67% and 51.95% for CA and GIS respectively. Interestingly, PS-Gurobi with random vectors occasionally surpasses the original model, likely due to the significant increase in input dimensionality - from 4 and 6 features for constraints and variables, respectively, to 68 and 70 - resulting in a larger model capacity that can yield marginal performance gains.
@@ 6 Related Work
Multi-Task and Generalist Models Efforts to unify learning across tasks and problem types have also emerged. Cai et al. (2025a) introduce a multi-task representation learning framework for MILP, training a shared backbone across tasks such as backdoor prediction and solver configuration prediction, followed by fine-tuning for specific problem types. Their method uses a similar bipartite graph representation, Graph Attention Networks (GAT), and contrastive loss, and is evaluated on problems like Combinatorial Auctions, Minimum Vertex Cover (MVC), and Maximum Independent Set (MIS). Similarly, Drakulic et al. (2024) present GOAL, a generalist agent for combinatorial optimization. It avoids GNNs, instead using mixed attention over edge and node matrices derived from bipartite graphs. Li et al. (2025) propose an LLM-based evolutionary framework that can generate a large set of diverse MILP classes and can be fine tuned to predict integrality gaps and branching nodes.
Graph Neural Networks for Branching and Heuristics Graph-based representations have become standard for encoding MILP instances. One of the earliest work by Gasse et al. (2019) uses GCNs to learn strong branching policies, introducing a bipartite graph structure and dual half-convolutions to facilitate message passing between constraints and variables. Chen et al. (2024) revisit GNN for MIPs and show that higher-order GNNs can overcome limitations identified via the 1-Weisfeiler-Lehman test, making all instances tractable for message passing. Cantürk et al. (2024) introduce improvements to the standard GNN workflow for CO problems so that they generalize on instances of a larger scale than those used in training and proposes a two-stage primal heuristic strategy based on uncertainty quantification to automatically configure how solution search relies on the predicted decision values.

### v2

@@ 5.2.1 Numerical Results: Integrality Gap Prediction
Test Instances & Setup: We start with the pre-trained Forge model, fine-tuned with a prediction head using the labeled data, and predict the integrality gap of 50 ‘very-hard’ instances each of CA, SC, GIS, and MVC. Given the prediction, a pseudo-cut is generated by adjusting the initial LP relaxation objective - by increasing it for minimization problems or decreasing it for maximization problems. This pseudo-cut is incorporated into the original formulation as an additional constraint, enforcing that the integral objective must exceed (or fall below) the computed pseudo-cut value.
Prediction Accuracy: We measure the deviation in terms of mean absolute error between the true integrality gap and the gap predicted by Forge. On the ‘very-hard’ instances, Forge achieves a deviation of $15.42\%$, $13.55\%$, $12.03\%$ and $19.077\%$ for CA, SC, GIS and MVC instances respectively. Note that MVC instances were not part of the train set.
@@
Comparison with the Commercial MIP Solver: We compare the default Gurobi solver on these ‘very-hard‘ instances with and without our predicted pseudo-cut. Figure 5 shows the primal gap averaged over these 50 ‘very-hard’ instances of each type with a time limit of 3600 seconds. Without exception, across all problem types, the use of pseudo-cuts generated by Forge embeddings consistently results in lower primal gaps over time. The gains range from 30% to 85%. The default solver attains immediate improvements early on in the search and our pseudo-cuts make these gains more pronounced immediately. Recall that this only needed 50 instances from each problem type.
@@
Comparison with SOTA ML Methods: To further evaluate generalization across problem types and sizes, we compare against the setup described in  Li et al. (2025), where a GNN is trained on 38,256 instances from 643 generated problem types and tested on 11,584 instances spanning 157 problem types. We use Forge as is, not using any additional training instances apart from 450 D-MIPLIB instances, and test on 17,500 previously unseen instances spanning 400 generated problem types. Forge achieves a mean deviation of 18.63% in integrality gap prediction, outperforming the reported 20.14% deviation by almost 10%.
@@ 5.3 Warm-Start Prediction & Guiding The Search
To address this issue, we generate 5 feasible solutions for each instance. Variables in the instance are grouped by how many of the 5 solutions each variable was repeated in. A variable is considered a “negative” variable if and only if it appears in none of the 5 solutions. Next, we use triplet loss (Schroff et al., 2015) to fine tune the embeddings of the variables. The triplet loss is given by
@@
Here, $a$ is the ‘anchor’ node, $p$ is the ‘positive’ node, $n$ is the ‘negative’ node and $d$ is a distance function (euclidean distance in our case). Triplet loss aims to minimize the distance between the ‘anchor’ and ‘positive’ nodes while ensuring the ‘negative’ node is at least ‘margin’ distance away from the ‘anchor’ node (margin is set to 2 in our case). In our setting, as shown in Figure 4(C), all variables appearing in the same number of solutions are treated as ‘positive’ and ‘anchor’ pairs. A key challenge of triplet loss is finding good negative nodes, as picking trivially negative nodes does not aid learning. We pick variables that have not appeared in any solution but are closest to the positive node in the unsupervised embedding space as negative nodes. Finally, the Forge model is fine-tuned to predict warm start variables using a combination of triplet loss and BCE loss.
Training Instances: We collect 100 instances each from CA (easy, medium), SC (easy, medium, hard) and GIS (easy, medium) for a total of 700 training instances.
Label Collection: Each instance is solved using Gurobi and a solution pool of 5. Here, we set a larger time limit of 300 seconds since we would like to get closer to the optimal solution.
@@ 5.3.1 Numerical Results: Warm-Start Prediction
Figure 7: Performance comparison of three variants of PS-Gurobi: the original model, a version augmented with 64-dimensional random vectors, and a version augmented with Forge embeddings (reduced to 64 dimensions via PCA). Evaluation is conducted on 50 ‘medium’ instances each of CA and GIS problems. The Forge-augmented PS-Gurobi consistently outperforms the baseline, achieving average reductions in primal gap of 41.07% for CA and 50.51% for GISP.
@@
Since Forge embeddings for each node are 1024-dimensional, we apply Principal Component Analysis (PCA) to reduce the dimensionality to 64 dimensions, thereby minimizing architectural modifications to the original model. To further ensure that any observed performance gains are attributed to the semantic content of the Forge embeddings rather than increased model capacity, we also conduct a control experiment by augmenting PS-Gurobi with 64-dimensional random vectors. This comparison allows us to isolate the contribution of the learned embeddings to overall performance. We evaluate this setup on the same 50 instances each of the CA and GIS problems - two problem types that are common to both Forge and PS-Gurobi. Results are shown in Figure 7 where PS-Gurobi + Forge consistently outperforms the standalone PS-Gurobi with the gain in primal gaps being 41.07% and 50.51% for CA and GIS respectively. Interestingly, PS-Gurobi with random vectors occasionally surpasses the original model, likely due to the significant increase in input dimensionality - from 4 and 6 features for constraints and variables, respectively, to 68 and 70 - resulting in a larger model capacity that can yield marginal performance gains.
@@ 6 Related Work
Multi-Task and Generalist Models Efforts to unify learning across tasks and problem types have also emerged. Cai et al. (2025a) introduce a multi-task representation learning framework for MIP, training a shared backbone across tasks such as backdoor prediction and solver configuration prediction, followed by fine-tuning for specific problem types. Their method uses a bipartite graph representation, Graph Attention Networks (GAT), and contrastive loss, and is evaluated on problems like CA, MVC and Maximum Independent Set (MIS). Similarly, Drakulic et al. (2024) present GOAL, a generalist agent for combinatorial optimization. It avoids GNNs, instead using mixed attention over edge and node matrices derived from bipartite graphs. Li et al. (2025) propose an LLM-based evolutionary framework that can generate a large set of diverse MILP classes and can be fine tuned to predict integrality gaps and branching nodes.
Graph Neural Networks for Branching and Heuristics Graph-based representations have become standard for encoding MIP instances. One of the earliest work by Gasse et al. (2019) uses GNNs to learn strong branching policies, introducing a bipartite graph structure and dual half-convolutions for message passing between constraints and variables. Chen et al. (2024) revisit GNN for MIPs and show that higher-order GNNs can overcome limitations identified via the 1-Weisfeiler-Lehman test, making all instances tractable for message passing. Cantürk et al. (2024) introduce improvements to the standard GNN workflow for CO so that they generalize on instances of a larger scale than those used in training and propose a two-stage primal heuristic strategy based on uncertainty quantification to automatically configure how solution search relies on the predicted decision values.